OpenStudy (anonymous):

Choose the graph below that correctly represents the equation 3x - 4y = 12. line through the points 0 comma 4 and 3 comma 0

OpenStudy (amoodarya):

put x=0 and find y then put y=0 and find x now you have two points line passes from to point s

OpenStudy (amoodarya):

3x - 4y = 12 x=0 0-4y=12 -4y=12 y=-3 so point is (0,-3) y=0 3x-0=12 3x=12 x=4 so point is (4,0)

OpenStudy (amoodarya):



OpenStudy (anonymous):

Thank you so much! :D